OpenStudy (anonymous):

Create your own two-step equation with All Real Numbers as the solution. Show the steps to solving the equation to prove the solution All Real Numbers.

OpenStudy (anonymous):

x^2+2x+1=0, is valid for all real x. so factorising, x^2 + x + x +1=0 x(x+1)+1(x+1)=0 (x+1)(x+1)=0 x+1=0 x=-1